# Minimally Entangled Typical Thermal States for Classical and Quantum Simulation of Gauge Theories at Finite Temperature and Density

### v1

@@ IV.4 System-size scaling of AVQMETTS circuit complexity
$N_{\textrm{CX}}=\sum_{j=1}^{M}2\left(W_{j}-1\right),$ (9)
where $W_{j}$ is the weight (number of nonidentity Pauli operators) of the Pauli string $G_{j}$.
@@
In Fig. 8(a) we show the statistics of $N_{\textrm{CX}}$ (symbols) for AVQITE circuits with initial CPSs in the $y$ and $z$ bases as a function of system size $L$. Here the confining field is set to $h=0$, and the chemical potential $\mu$ is adjusted such that the total fermion filling $\langle N\rangle=L/4$. We apply AVQITE to propagate the CPSs up to a temperature $\beta=1,2$. The dashed lines show that they can be fitted well by a simple power-law function $f(L)=aL^{b}$, with exponent $b$ ranging from $1.6$ to $2.1$ for the different measurement basis and temperature values considered here. One can clearly see a bimodal distribution of AVQITE calculations: $N_{\textrm{CX}}$ for AVQITE calculations starting from $z$-basis CPSs is in the range of $[4,30]$, while that of $y$-basis in $[20,800]$, demonstrating over an order of magnitude increase at fixed $L$. Furthermore, increasing $\beta$ at fixed $L$ leads to only a slight increase in $N_{\textrm{CX}}$ for $z$-basis results, in contrast to the substantially larger increase observed for the $y$-basis results. Fig. 8 shows that $N_{\textrm{CX}}$ displays a modest up-shift as we change the confining field from $h=0$ to $h=0.1$; the maximal $N_{\textrm{CX}}$ rises from $32$ to $34$ for the $z$-basis results, and from $740$ to $900$ for $y$-basis results. Hence, even though we showed in Table 1 that the $y$-basis samples are more concentrated around the thermal average, the implied greater sampling efficiency can be compromised by the increased quantum resource overhead in AVQMETTS calculations. This again suggests that AVQMETTS using the alternating $yz$-basis for state collapse may be an optimal choice.
@@ V Conclusion
The methods adopted in this work to calculate the equation of state may offer insight into two topics of current interest in the study of neutron matter, namely neutron star mass-radius relations [56, 57, 58, 59] and multi-messenger signals of neutron star mergers [60, 61, 62, 63, 64]. Both require a knowledge of the equation of state of nuclear matter at finite density and temperature. In the case of the former, neutron stars data which favors a soft equation of state at low density and a stiff one at high density is in tension with several theoretical models which predict a rather soft equation of state at high density [65]. In fact this soft-stiff equation of state data from neutron stars suggests that the speed of sound is non-monotonic as a function of density. Since first-principles QCD calculations are currently out of reach, it is worthwhile exploring whether toy models in lower dimensional theories like the one studied in this work can exhibit some of these features expected in the QCD equation of state. Similarly, nuclear/QCD equations of state at relatively higher temperature are used in numerical relativity simulations to predict gravitational wave signals from neutron star merger events. Computing finite temperature equations of state for toy models is a step towards eventually doing the same for QCD. By investigating the systematics of (AVQ)METTS calculations, our paper lays the groundwork for both classical and quantum computers to tackle these questions in the future.

### v2

@@ IV.4 System-size scaling of AVQMETTS circuit complexity
$N_{\textrm{CX}}=\sum_{j=1}^{N_{\theta}}2\left(W_{j}-1\right),$ (9)
where $W_{j}$ is the weight (number of nonidentity Pauli operators) of the Pauli string $G_{j}$. Notice that Eq. (9) sets an upper bound for the total number of CNOT gates: it does not account for possible gate cancellations in the translation of the ansatz to a quantum circuit.
@@
In Fig. 8(a) we show the statistics of $N_{\textrm{CX}}$ (symbols) for AVQITE circuits with initial CPSs in the $y$ and $z$ bases as a function of system size $L$. Here the confining field is set to $h=0$, and the chemical potential $\mu$ is adjusted such that the total fermion filling $\langle N\rangle=L/4$. We apply AVQITE to propagate the CPSs up to temperatures $\beta=1,2$. The dashed lines show that the data are fitted well by a simple power-law function $f(L)=aL^{b}$, with exponent $b$ ranging from $1.6$ to $2.1$ for the different measurement basis and temperature values considered here. One can clearly see a bimodal distribution of AVQITE calculations: for calculations starting from $z$-basis CPSs, $N_{\textrm{CX}}$ is in the range of $[4,30]$, while for ones starting from $y$-basis CPSs it is in the range $[20,800]$, demonstrating over an order of magnitude increase at fixed $L$. Furthermore, increasing $\beta$ at fixed $L$ leads to only a slight increase in $N_{\textrm{CX}}$ for $z$-basis results, in contrast to the substantially larger increase observed for the $y$-basis results. Hence, even though we showed in Table 1 that the $y$-basis samples are more concentrated around the thermal average, the implied greater sampling efficiency can be compromised by the increased quantum resource overhead in AVQMETTS calculations. This again suggests that AVQMETTS using the alternating $yz$-basis for state collapse may be an optimal choice.
In Fig. 8(b) and (c) we extend the results for $N_{\textrm{CX}}$ to larger system sizes (up to $L=32$) using a tensor-network (TN) simulator of the AVQITE algorithm [58] based on the Quimb [59] and cotengra packages [60]. Unlike the statevector simulations discussed above, where we sampled over $288$ initial CPSs, here we instead sample over $30$ initial CPSs. Nevertheless, we have checked that doubling the number of samples to $60$ does not meaningfully alter the results. In addition, due to the higher circuit depths for the $y$-basis samples, we perform TN simulations for the initial CPSs in the $z$-basis only. However, there is no obstruction to performing the $y$-basis calculations, increasing the number of samples, and considering system sizes larger than $L=32$ given sufficient HPC resources.
Fig. 8(b) demonstrates that for the larger system sizes, the means of the $N_{\textrm{CX}}$ distributions still follow the power-law dependence $f(L)=aL^{b}$, with the exponent $b$ ranging between $1.6$ and $2.1$. Additionally, Fig. 8(c) shows that the power-law dependence of $N_{\textrm{CX}}$ persists as we change the confining field from $h=0$ to $h=0.1$. Note that the mismatch in $N_{\textrm{CX}}$ between Fig. 8(a) and (b) at $L=8$ and $L=16$ comes from the fact that in (a) the implementation of AVQITE prioritizes the selection of low-weight generators if their scores are close, whereas in (b) this is not the case and the generators with close scores are sorted in alphabetical order. We, however, do not expect such fine tunings of the algorithm to affect the scaling behavior of $N_{\textrm{CX}}$, and indeed panels (a) and (b) yield consistent estimates of this scaling.
These results showcase that TN-based simulations can serve as a promising avenue for benchmarking and analyzing the system-size scaling complexity of quantum algorithms for system sizes beyond the reach of statevector simulators. This is especially relevant for heuristic quantum algorithms (such as variational ones), for which scaling behavior and use-case-appropriate bounds on other resource requirements generally cannot be obtained through analytical arguments [61].
In addition, the TN simulator can be leveraged by combining it with a QPU in a hybrid quantum-classical workflow for the AVQITE algorithm. In such a workflow, the TN approach would be used to classically simulate quantum circuits that are amenable to such simulations (e.g., circuits with shallower ansätze that comprise a block of the quantum geometric tensor corresponding to the early stages of the algorithm), while a QPU would only be utilized for circuits whose complexity is inaccessible to state of the art TN techniques. Crucially, for any given quantum circuit, the computational and memory requirements of its TN contraction (i.e. the contraction cost and width) can be estimated prior to evaluating the contraction, from which the feasibility of such classical evaluation can be assessed. The described hybrid TN-QPU workflow can considerably reduce the number of QPU circuit evaluations needed to scale AVQITE to the beyond-classical regime [62].
IV.5 Alternative Methods
Having presented a detailed benchmark of the AVQMETTS approach to simulating finite-temperature and -density properties of the 1+1 dimensional $\mathbb{Z}_{2}$ lattice gauge theory, we briefly survey some alternative approaches to the problem. In general, simulation of finite temperature physics on a quantum computer is more involved than zero-temperature as quantum computers do not naturally operate on mixed states.
The METTS approach explored in this work can be used in combination with any algorithm implementing the imaginary time evolution primitive on a quantum computer. Besides the variational approach adopted here, several other implementations of the ITE primitive have been proposed. Ref. [43] proposes to perform ITE with the help of state tomography on a neighborhood whose size scales with the correlation length. ITE can also be implemented using a register of ancilla qubits as well as mid-circuit measurements or resets, as in Refs. [63, 64]. More broadly, ITE can be implemented using generic primitives such as quantum signal processing [65, 66, 67, 68, 69] or block encoding [70, 71, 72, 73, 74].
There are also alternative approaches to estimating thermal averages on quantum computers. Some of these assume access to ITE as a primitive. The “typicality” based methods can compute finite-temperature averages by evolving a fully random or pseudo-random state [69, 31] of the entire system with the preparation of this pseudo-random state being efficient in principle. Another approach gives access to a purification of the full thermal density matrix by performing imaginary time evolution on half of a collection of Bell pairs [55]. Finally, one can also consider fully variational methods for preparing mixed states by parameterizing $\rho(\boldsymbol{\theta},\boldsymbol{\phi})=\sum_{k}p_{k}(\boldsymbol{\theta})\ket{\psi_{k}(\boldsymbol{\phi})}\bra{\psi_{k}(\boldsymbol{\phi})}$ | and minimizing the free energy [75]. Due to the difficulty of measuring the entropic term, other variants of this scheme have been proposed using, for example, mid-circuit measurements [76]. Microcanonical averages may also be accessible by generating variational samples of the microcanonical ensemble [77].
As with AVQMETTS, the scalability of these various approaches in practice for near-term hardware is not yet established. It will be necessary to continue investigating means to reduce obstacles such as measurement, qubit, and gate overheads for all of these approaches. Furthermore, these methods are all impacted in different ways by gate errors, measurement errors, and shot noise. As hardware and algorithms continue to evolve, careful benchmarks like the ones performed in this paper are necessary to gauge the feasibility of finite-temperature quantum simulation on quantum computers.
@@ V Conclusion
Although we have focused in this paper on the simplest 1+1-dimensional $\mathbb{Z}_{2}$ lattice gauge theory, we note that many of the techniques adopted here can potentially be applied to higher-dimensional systems and for other gauge groups. For example, a 2+1 dimensional version of the $\mathbb{Z}_{2}$ lattice gauge theory studied in this work has recently been shown to map to a two-dimensional qubit-only model [78]. Spin-1/2 formulations of $U(1)$ lattice gauge theory in 1+1 and 2+1 dimensions [26, 79] are amenable to digital quantum simulation with qubits, and some initial progress has been made in understanding the confining dynamics in the 1+1 dimensional case [80]. The 2+1 dimensional models can be studied using METTS on narrow cylinders, as has recently been done for some models of correlated electron systems [41, 42]. There are also prospects for simulating non-Abelian gauge theories in 1+1 [81, 82] and 2+1 dimensions [83] using tensor networks, and for quantum simulations of such models in, e.g., trapped-ion quantum computers [84]. More generally, encodings into qubit- or qudit-only models have been proposed for U($N$) and SU($N$) gauge theories with fermionic matter in 2+1 dimensions [85]. For the simplest of these models (e.g. the Abelian ones), quantum simulation of finite-temperature properties using the methods explored in this work may soon be feasible. For the more complicated models with tensor-network formulations, finite-temperature classical simulations using METTS are a natural next step.
The methods adopted in this work to calculate the equation of state may offer insight into two topics of current interest in the study of neutron matter, namely neutron star mass-radius relations [86, 87, 88, 89] and multi-messenger signals of neutron star mergers [90, 91, 92, 93, 94]. Both require a knowledge of the equation of state of nuclear matter at finite density and temperature. In the case of the former, neutron star observational data which favor a soft equation of state at low density and a stiff one at high density are in tension with several theoretical models which predict a rather soft equation of state at high density [95]. In fact this soft-stiff equation of state data from neutron stars suggests that the speed of sound is non-monotonic as a function of density. Since first-principles QCD calculations are currently out of reach, it is worthwhile exploring whether toy models in lower dimensional theories like the one studied in this work can exhibit some of these features expected in the QCD equation of state. Similarly, nuclear/QCD equations of state at relatively higher temperature are used in numerical relativity simulations to predict gravitational wave signals from neutron star merger events. Computing finite temperature equations of state for toy models is a step towards eventually doing the same for QCD. By investigating the systematics of (AVQ)METTS calculations, our paper lays the groundwork for both classical and quantum computers to tackle these questions in the future.
Data availability
Data used in all plots in this manuscript are available at https://gitlab.com/gqce/avqmetts_z2.
Code availability
Code used to produce all plots in this manuscript is available at https://gitlab.com/gqce/avqmetts_z2.